zdfbzdfbzdfb:

Joe purchased a tray of 48 petunias to plant in his garden. The graph shows points representing the number of flowers left to plant at several different times while Joe is planting. Select from the drop-down menus to complete the equation, in standard form, that represents the relationship between the number of petunias left to plant, y, and the time, x, Joe has spent planting. Choose... x + y = Choose... A

ihatelifeFS420:

do you have answer choices or no

zdfbzdfbzdfb:

i can give you them yes

zdfbzdfbzdfb:

choose... 1 12 36 48 choose... 1 4 12 48

zdfbzdfbzdfb:

1 attachment
ihatelifeFS420:

so you have to find the slope=(48-36)/ (0,-1)=-12 then it can be written as y=-12x+48 or 12x+y=48

zdfbzdfbzdfb:

ty

ihatelifeFS420:

@zdfbzdfbzdfb wrote:
ty
yw